OpenStudy (anonymous):

In a solution being diluted, the amount of solute remains constant. How many liters of 8.0 molar Ca(0H)2 stock solution will be needed to make 5.0 liters of a 1.5 molar Ca(0H)2 solution? Use M(concentration) * Volume(concentration) = M(dilution) * V(dilution)

OpenStudy (anonymous):

If someone could please do this for me, thank you so much!!

OpenStudy (anonymous):

Show all calculations

OpenStudy (anonymous):

Btw, M= moles, V= volume

OpenStudy (abb0t):

Why can't you do it? It's pretty straight forward. You have \(\sf \color{red}{M_1V_1=M_2V_2}\) and you solve for \(\sf \color{blue}{ V_1}\) You are given: M\(_1\) = 8.0 m V\(_2\) = 5.0 L M\(_2\) = 1.5 M

OpenStudy (abb0t):

Note that stock solution means your original solution. What you started with to essentially dilute to down to 1.5 M

OpenStudy (anonymous):

I just don't get it

OpenStudy (abb0t):

by the way, M ≠ moles M is actually \(\sf \color{Red}{molarity}\) which is moles of solute per liter of solution, \(\sf \frac{moles}{Liter}\)

OpenStudy (abb0t):

Yes, you can do it. Just try it. I already told you what you have. You have \(\sf \color{red}{M_1V_1=M_2V_2}\) Think back to basic algebra, how would you solve for \(\sf \color{blue}{V_1}\)? Rearranging the formula...

OpenStudy (anonymous):

Do I just plug everything in

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so would the M1 go to the right side like this? V1= M2V2M1

OpenStudy (abb0t):

Yes. That is all you do. But, you have to kind of tweek the formula a bit, since you're solving for V\(_1\). Since it is asking you "how many liters of stock solution", which like i said before is your initial

OpenStudy (abb0t):

You're close, but not quite. If you are multiplying, you do the opposite. That means, you DIVIDE by M\(_1\). Not multiply.

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

V1= M2*V2/M1

OpenStudy (abb0t):

Yes. Correct: \(\sf \color{BLUE}{V_1}=\frac{M_2V_2}{\color{RED}{M_1}}\)

OpenStudy (anonymous):

can you help me with the rest

OpenStudy (abb0t):

That's all. You're done. Just plug it and solve. Use a calculator.

OpenStudy (anonymous):

ok